Determine the unknown currents using the current d
Answer #1

Here,

for the resistance of 8 and 12 Ohm in parallel

R1 = 8 * 12/(8 + 12)

R1 = 4.8 Ohm

Using current divider

I1 = 6 * 4.8/(4.8 + 4)

I1 = 3.27 A

again

I2 = (6 - 3.27) * 12/(12 + 8)

I2 = 1.64 Ohm

again

I3 = (6 - 3.27) * 8/(12 + 8)

I3 = 1.09 A

and current , I4 = 6 A